OpenStudy (anonymous):

Evaluate \[\lim_{x \rightarrow \frac{ \pi^- }{ 2 }}e ^{tanx}\]

OpenStudy (anonymous):

I don't understand why the answer is positive infinity.

OpenStudy (misty1212):

\[\huge \lim_{x \rightarrow \frac{ \pi}{ 2 }^-}e ^{\tan(x)}\]

OpenStudy (rational):

\(e^x\) is a continuous function, so \[\lim (e^{f(x)}) = e^{\lim (f(x))}\]

OpenStudy (misty1212):

just trying to read it is all

OpenStudy (misty1212):

goes to \(\frac{\pi}{2}\) from the left

OpenStudy (misty1212):



OpenStudy (misty1212):

tangent is going to \(+\infty\)

OpenStudy (anonymous):

what @misty1212 said tangent goes to plus infinity so \(e^{\tan(x)}\) goes to infinity real real real fast

OpenStudy (anonymous):

Oh okay. Just wondering can I sub in the pi/2 into the tanx? Does that help me?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

since \(\frac{\pi}{2}\) is not in the domain of tangent you cannot do that

OpenStudy (anonymous):

otherwise that would be the right thing to do, but not in this case

OpenStudy (anonymous):

So the only way to solve this is to think logically?

OpenStudy (anonymous):

heaven forbid

OpenStudy (anonymous):

i mean "yes"

OpenStudy (rational):



OpenStudy (anonymous):

oh, pi/2 is vertical asymptote

OpenStudy (rational):

its much worse than that, as you walk from left side it goes to +infinity as you walk from right side it goes to -infinity

OpenStudy (rational):

so sometimes we say the two sided limit for \(\tan x\) as \(x\to \pi/2\) does not exist

OpenStudy (anonymous):

Okay, thank you!